OpenStudy (yacoub1993):

Can someone explain to me step by step.. Verify the identity. cot θ ∙ sec θ = csc θ

OpenStudy (anonymous):

\[\cos \theta \div \sin \theta \times 1\div \cos \theta = 1\div \sin \theta \] then the cos cancels out, and you end up with 1/sin= 1/sin (hope that helps!)